# Know the Unknown: An Uncertainty-Sensitive Method for LLM Instruction Tuning

### v1

@@ 4 Experiments
Model $Acc_{known}$ $Acc_{unknown}$ $F1$ $Acc_{known}$ $Acc_{unknown}$ $F1$
GPT-4 Turbo 79.6 83.6 81.6 88.9 78.3 83.3
@@
Llama2-Chat-7B (base) 79.3 58.3 67.2 85.2 29.5 43.9
@@
Table 2: Results (in %) for mainstream and our finetuned LLMs on QA uncertainty-recognition assessment and postfix uncertainty-recognition assessment. The overall best results are highlighted in bold. The results obviously higher and lower than the baseline are highlighted in green and orange, respectively.
4.1 Data Construction
Uncertainty-Sensitive Tuning Dataset In uncertainty-sensitive tuning, there are two distinct types of instruction datasets: the binary uncertainty-recognition tuning dataset and the prompt-sensitive fine-tuning dataset. For the binary uncertainty-recognition tuning dataset, we sample 646 examples from the ASQA dataset Stelmakh et al. (2022), which consists of ambiguous question-context pairs. Each pair is divided into multiple sub-questions with their corresponding contexts. For example, one pair may discuss the discovery of the photoelectric effect in 1887, while another pair may cover the theoretical development of the photoelectric effect in 1905. To generate the unknown questions, we shuffle these pairs, reassigning the questions to different but related contexts. The ambiguity between question-context pairs introduces minor deviations that challenge the model’s ability to handle uncertainty, differentiating our method from other works. The prompt-sensitive fine-tuning dataset is based on HotpotQA Yang et al. (2018), a dataset designed for multi-hop reasoning question answering. To adapt this dataset for fine-tuning with additional causal instructions, we generate and manually select the most effective 28 causal instructions. Then, we integrate these instructions with 300 randomly selected samples from HotpotQA. Finally, all the generated datasets are used for fine-tuning. Here is a demonstration of the additional prompt in uncertainty-recognition tuning:
Uncertainty-Recognition Benchmark To comprehensively evaluate the model’s cognitive ability to recognize knowledge gaps, we construct a testbed using the ASQA Stelmakh et al. (2022)dataset, ensuring no overlap with the training data. During the evaluation, two types of designed evaluation templates are available: the QA uncertainty recognition template and the postfix uncertainty recognition template. For the QA uncertainty recognition, we instruct the model to either answer the question or reject it with ’Not Provided’. We generate 3320 test data for this type of template. For postfix uncertainty recognition, we instruct the LLMs to answer the question before classifying it. We include 646 test samples for this template. Here is a demonstration of the used prompt templates:
@@ 4.2 Experiment Setting
Training Details. Our experiments are conducted on mainstream LLMs, including Llama2-Chat-7B Touvron et al. (2023), GPT-4 Turbo OpenAI (2023) , GPT-3.5 Turbo OpenAI (2023), Vicuna-7B v1.5 Zheng et al. (2024a) and Self-RAG-7B Asai et al. (2023). Taking Llama2 as the baseline, we further finetune it on a single GTX3090 GPU by LLaMA-Factory Zheng et al. (2024b). Training with Lora Hu et al. (2021) in a rank of 8, with a batch size of 4 and a learning rate of 5e-5. This research integrates the instruction-based prompt and attributed prompt which is proved to be effective for mitigating hallucination Zhou et al. (2023), as provided in Appx. A. We fine-tuned three different versions of the model, each using distinct instruction datasets. The uncertainty-recognition-tuned model was fine-tuned using the proposed uncertainty-recognition-tuning dataset. The Uncertainty & HotpotQA tuned model was fine-tuned with both the proposed uncertainty-recognition tuning dataset and the HotpotQA dataset. Lastly, the uncertainty-sensitivity tuned model utilized a combination of the uncertainty-recognition tuning dataset and the prompt-sensitivity tuning dataset.
Evaluation Metric. We use $Acc_{known}$ to demonstrate the accuracy of questions with specific answers and $Acc_{unknown}$ represents the accuracy of unknown questions.
@@ 4.3 Results
Table 2 presents the results from our evaluation of mainstream large language models (LLMs). The evaluation of question-answering (QA) uncertainty reveals that most large language models (LLMs) struggle to accurately identify unknown questions, with only around 50% accuracy. However, GPT-4 stands out as a notable exception, outperforming the second-best model, Llama2, by a significant margin of 25.3%. Llama2 ranks highest among the remaining models, even surpassing GPT-3.5, despite the latter having more parameters. But there is still a substantial performance gap between the GPT-4 and the other models. The ongoing experiments aim to investigate the reasons behind this disparity. The postfix uncertainty-recognition assessment results, shown in the right column of Table 2, indicate a strong positive correlation between the two evaluation templates. The analysis also highlights that different models respond differently to insufficient information. Models fine-tuned on dialogue tasks tend to overly rely on and trust the given information. Self-RAG, which is fine-tuned for question-answering tasks with unknown questions, demonstrates a strong ability to reject questions it cannot answer. However, despite Self-RAG’s impressive performance in the uncertainty-recognition assessment, it remains difficult to train the model to effectively acknowledge unknown information.
Our fine-tuned model, which has been tuned for uncertainty recognition, performs comparably to GPT-4 in tests of uncertainty recognition and shows a 25.9% improvement in responding to context compared to the baseline model. However, despite outperforming GPT-4 in terms of unknown questions ($Acc_{unknown}$), this increased awareness of uncertainty leads to a decrease in the model’s original question-answering capabilities. Further fine-tuning on the HotpotQA dataset resulted in the model being less reliable in following instructions and sometimes disregarding its knowledge limitations. Among the models tested, the uncertainty-sensitivity tuned model ranks the highest, achieving an F1 score of 85.8, which is 4.2 points higher than GPT-4 and 18.6 points above the baseline (as shown in Figure 2). This impressive performance is due to the model’s ideal balance between answering questions and acknowledging when it lacks the necessary information to provide an answer. It achieves a remarkable 93.0% accuracy on unknown questions, the highest among all mainstream LLMs, while maintaining a 79.7% accuracy on known questions. Moreover, the model effectively integrates new information without compromising its existing knowledge base, perfectly retaining the knowledge it acquired from cognition fine-tuning.
@@ 4.4.1 Experiment Setting
In order to further quantify the effects of uncertainty-recognition tuning, HotpotQA tuning and prompt-sensitive tuning, three additional variants have been built for ablation experiments.
  • The HotpotQA finetuning utilizes 300 samples with traditional question answering prompts provided in Appx. A.1.
@@
  • The Uncertainty & HotpotQA tuned model further finetunes the uncertainty-recognition tuned model on HotpotQA dataset.
@@ 4.4.2 Result Analysis
Weak Uncertainty-Recognition Capacity: Our ablation study, as shown in Fig. 3, illustrates that enhancing the model’s ability to recognize uncertainty significantly boosts performance on unknown questions. Specifically, the uncertainty-recognition tuning resulted in a marked 16.9% increase in ${Acc}_{unknown}$. The models that have not been trained on unknown tasks exhibit poor performance. We call this weak uncertainty recognition. The problem of weak uncertainty-recognition capacity in LLMs comes from their fundamental training process to generate specific responses. In scenarios where LLMs encounter questions outside their training distribution or scope of knowledge, they often fail to adequately recognize uncertainty. This leads to inappropriate or misleading answers.
Prompt-sensitivity reduction problem: The reduction in known questions is another critical issue highlighted by the ablation study. As models are further tuned to handle specific tasks like QA finetuning or uncertainty recognition, they can become less responsive to variations in the input prompt, a phenomenon we term "prompt-sensitivity reduction." This reduced sensitivity can impair the model’s ability to adapt to the slight but significant differences in how questions are phrased or presented, leading to what we describe as cognitive inaccessibility. From Fig. 3, it is evident that although simple QA finetuning attempts to mitigate the negative effects of over-tuning for uncertainty recognition, it does not fully address the prompt-sensitivity issue. The model continues to struggle with lower overall accuracy because it cannot effectively adapt to the diverse ways in which users might phrase similar questions. Thus, the model fails to fully leverage the context of the prompt, resulting in inaccurate or relevant answers. With our prompt-sensitive tuning, we can successfully recover the performance of other questions. Even if we do not fine-tune the model for the unknown task, our prompt-sensitive tuning can boost the performance of the task with specific answers, which is the highest accuracy in the table.
@@
Llama2-Chat-7B (baseline) 79.3 58.3 67.2
@@
Figure 3: Comparison of different models with different training dataset. A higher score means better performance.
@@ 5 Conclusion
In this paper, we explore a common issue in language model applications where insufficient context in prompts leads to the generation of plausible yet incorrect responses, a phenomenon known as hallucinations. Firstly, our research identifies that large language models (LLMs) often fail to recognize when they lack adequate information to accurately respond to a query. We primarily attribute this problem to the fact that the models lack the ability to recognize when the provided context is insufficient for the posed queries. Second, during the training process with questions without a specific answer, the performance of questions with answers decreased. To counter these issues, we introduce a novel training methodology termed "uncertainty-sensitive tuning."  This is a two-stage training framework. In the first stage, we guide the LLM to reject unknown questions. In the second stage, we recover the decreased performance in QA tasks by incorporating designed causal instructions. This approach significantly enhances the Llama2-chat-7B model’s ability to handle queries more reliably, reducing the occurrence of hallucinations. Our approach differs from previous works in that it focuses more directly on refining the model’s response behavior through targeted fine-tuning, thereby enhancing its performance in practical scenarios. By open-source this framework and model, we believe that it will highlight a new direction for the automatic synthesis instruction data set, which is not only focused on the diversity of the data but also the reduction of hallucinations during training.
@@ A.4 Prompt-Sensitive Finetuning
You should check whether your answer aligned the requirement by generating a Checking part, checking each sentence of the above instruction, with either <fulfilled> or <not fulfilled> mark behind the sentence, indicating whether the requirement is fulfilled or not. If there is <not fulfilled> mark behind the sentence, you must modify your answer again to fulfill the requirement, by appending a new ANSWER and CHECKING part. Here is an example for this task: e.g. Question Answering Task Requirements: You need to do the Task Prompt for the following query and context. Ensure the response is written in the past tense. QUESTION: Who is Jack Chen? CONTEXTS: People saying that Jack Chen is a famous singer in China. ANSWER: Jack Chen was a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<not fulfilled> ANSWER: Jack Chen was a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<fulfilled> Here is the information of your task: {Question Answering Instruction}

### v2

@@ 4.1 Data Construction
Category Model $Acc_{known}$ $Acc_{unknown}$ $F1$ $Acc_{known}$ $Acc_{unknown}$ $F1$
Benchmark GPT-4 Turbo 79.6 83.6 81.6 88.9 78.3 83.3
@@
Llama2 Llama2-7B-Chat(baseline) 79.3 58.3 67.2 85.2 29.5 43.9
@@
Mistral Mistral-7B-Instruct-v0.2(baseline) 85.1 63.0 72.4 82.8 43.1 56.7
Uncertainty-Recognition Tuned 77.5 75.8 76.6 86.1 81.9 84.0
Uncertainty & HotpotQA Tuned 87.1 52.4 65.5 80.7 75.1 77.8
Uncertainty-Sensitivity Tuned 87.3 75.3 80.9 82.5 82.2 82.4
Gemma Gemma-2-9B-Instruct(baseline) 86.1 74.1 73.5 86.3 57.6 69.1
Uncertainty-Recognition Tuned 76.1 86.2 80.8 93.4 76.2 83.9
Uncertainty & HotpotQA Tuned 91.3 20.8 33.9 99.4 58.7 73.8
Uncertainty-Sensitivity Tuned 87.6 81.2 84.3 96.1 55.1 70.1
Table 2: Results (in %) for mainstream and our finetuned LLMs on QA uncertainty-recognition assessment and postfix uncertainty-recognition assessment. The overall best results are highlighted in<textbfbold. The results obviously higher and lower than the baseline are highlighted in green and orange, respectively.
@@ 4.2 Experiment Setting
Training Details. We experiment our uncertainty-sensitive tuning approach on mainstream open-sourced LLMs, including Llama2-7B-Chat Touvron et al. (2023), Mistral-7B-Instruct-v0.2 Jiang et al. (2023), and Gemma-2-9B-Instruct Team et al. (2024). We also test GPT-4 Turbo OpenAI (2023), GPT-3.5 Turbo OpenAI (2023), Vicuna-7B v1.5 Zheng et al. (2024a) and Self-RAG-7B Asai et al. (2023) on our benchmark. Our finetuning bases on a single RTX3090 GPU by LLaMA-Factory Zheng et al. (2024b), with Lora Hu et al. (2021) in a rank of 8, with a batch size of 4 and a learning rate of 5e-5. This research integrates the instruction-based and attributed prompt which is proved to be effective for mitigating hallucination Zhou et al. (2023), as provided in Appx. A.
Evaluation Metric. We use $Acc_{known}$ for representing the accuracy of questions with specific answers, and $Acc_{unknown}$ for unknown questions.
@@ 4.3 Results
Table 2 presents the results of our benchmark. The evaluation of QA uncertainty-recognition reveals that most mainstream LLMs struggle to accurately identify unknown questions, with only around 60% accuracy. However, GPT-4 and Gemma-2 are notable exceptions with accuracies of 83.6% and 74.1%, respectively. Mistral and Llama-2 rank highest among the remaining models, even surpassing GPT-3.5, despite the latter having more parameters. But there is still a substantial performance gap between the GPT-4 and the other models. The ongoing experiments aim to investigate the reasons behind this disparity. The postfix uncertainty-recognition assessment results, shown in the right column of Table 2, indicate a strong positive correlation between the two evaluation templates. The analysis also highlights that different models respond differently to insufficient queries. Models fine-tuned on dialogue tasks tend to overly rely on and trust the given information. Self-RAG, which is fine-tuned for QA tasks with unknown questions, demonstrates a strong ability to identify the questions it cannot answer. However, it remains difficult for the model to admit the lack of knowledge.
Our uncertainty-recognition tuned models perform comparably to GPT-4 in tests of uncertainty recognition, and show an improvement up to 25.9% in responding to unknown questions compared to the baselines. However, despite outperforming GPT-4 in terms of unknown questions ($Acc_{unknown}$), this increased awareness of uncertainty leads to a decrease in original QA capabilities. Models are overly sensitive to the diverse ways in which users might phrase similar questions. Further fine-tuning on the HotpotQA resulted in the model being less reliable in following instructions and sometimes disregarding its knowledge limitations, despite the advancements in postfix uncertainty-recognition indicates the knowledge is maintained. Among the models tested, the uncertainty-sensitivity tuned Llama2 ranks the highest, achieving an F1 score of 85.8, which is 4.2 points higher than GPT-4 and 18.6 points above the baseline (as shown in Figure 2). This impressive performance is due to the model’s ideal balance between QA and rejecting answering. It achieves a remarkable 93.0% accuracy on unknown questions, the highest among all mainstream LLMs, while maintaining a 79.7% accuracy on known questions. Gemma-2 and Mistral also achieve 10.8% and 8.5% improvements in overall performance, demonstrating that our uncertainty-sensitivity tuning approach allows models to effectively integrate new information without compromising the knowledge acquired from uncertainty-recognition fine-tuning.
@@ 4.4.1 Experiment Setting
To further quantify the effects of uncertainty-recognition tuning, HotpotQA tuning and prompt-sensitive tuning, three additional variants have been built on Llama2-7B-Chat for ablation experiments.
  • The HotpotQA fine-tuning utilizes 300 samples with traditional question answering prompts provided in Appx. A.1.
@@
  • The Uncertainty & HotpotQA tuned model further fine-tunes the uncertainty-recognition tuned model on HotpotQA dataset.
@@ 4.4.2 Result Analysis
Weak Uncertainty-Recognition Capacity: Our ablation study, as shown in Table 3, illustrates that enhancing the model’s recognize-uncertainty ability significantly boosts performance on unknown questions. Specifically, the uncertainty-recognition tuning resulted in a marked 16.9% increase in ${Acc}_{unknown}$. The problem of weak uncertainty-recognition capacity in LLMs comes from their fundamental training process to generate specific responses. In scenarios where LLMs encounter questions outside their training distribution or scope of knowledge, they often fail to adequately recognize uncertainty. This leads to inappropriate or misleading answers.
@@
Llama2-7B-Chat (baseline) 79.3 58.3 67.2
@@
Figure 3: Comparison of different models in ablation study. A higher score means better performance.
Prompt-sensitivity reduction problem: The reduction in known questions is another critical issue highlighted by the ablation study. As models are further tuned on specific tasks, such as QA or uncertainty recognition, they can become less responsive to variations in the input prompt, a phenomenon we term "prompt-sensitivity reduction." This reduced sensitivity can impair the model’s ability to adapt to the slight but significant differences in how questions are phrased or presented, leading to what we describe as cognitive inaccessibility. From Table 3, it is evident that although tuning on HotpotQA attempts to mitigate the negative effects of over-tuning for uncertainty-recognition, it does not fully address the prompt-sensitivity issue. The model even ignores the constraints in prompts and continues to hallucinate answers for unknown questions. Thus, the model fails to fully leverage the context of the prompt, resulting in inaccurate or relevant answers. With our prompt-sensitive tuning, we can successfully recover the performance of other questions. Even if we do not fine-tune the model for the unknown task, our prompt-sensitive tuning can boost the performance of the task with specific answers, which is the highest accuracy in the table.
4.5 Comparisons with Post-Generation Methods
Cole et al. (2023) and Kadavath et al. (2022) propose using sampling and LM-validation for post-generation uncertainty detection. Here we compare their approaches with uncertainty-sensitive tuning. For sampling method, we sample three times from the Llama2-7B-Chat with 0.6 temperature, and adopt the most repeated answers. In LM-Validation we allow the model to additionally edit the generated answers. As recorded in Table 4, due to the weak uncertainty-recognition capacity, models struggle to identify the unknown questions.
Model $Acc_{pos}$ $Acc_{neg}$ $F1$
Llama2-7B-Chat (baseline) 79.3 58.3 67.2
Multi-Sampling 79.7 66.5 72.5
LM-Validation 82.5 53.8 65.1
Uncertainty-Sensitivity Tuned 79.7 93.0 85.8
Table 4: Comparison results with the post-generation methods on QA uncertainty-recognition benchmark.
@@ 5 Conclusion
In this paper, we explore a common issue in large language models (LLMs) where insufficient context in prompts leads to the generation of plausible yet incorrect responses. Firstly, our research identifies that LLMs often fail to recognize when they lack adequate information to accurately respond. We primarily attribute this problem to the models’ weak uncertainty-recognition capacity and overconfidence in their generated potential answers. Second, training on the questions with specific answers significantly decreases the performance of questions without answers. To counter these issues, we introduce a novel two-stage training framework termed "uncertainty-sensitive tuning". In the first stage, we guide the LLM to identify unknown questions. In the second stage, we recover the decreased QA performance by incorporating designed causal instructions. This approach significantly enhances the model’s reliability, reducing the occurrence of hallucinations. Our approach differs from previous works in that it focuses more directly on refining the response behavior through targeted fine-tuning, thereby enhancing its performance in practice. By open-sourcing this framework and dataset, we believe that it will highlight a new direction for the automatic synthesis instruction dataset, which is not only focused on the diversity of the data but also the reduction of hallucinations during training.
@@ A.4 Prompt-Sensitive Finetuning
You should check whether your answer aligned the requirement by generating a Checking part, checking each sentence of the above instruction, with either <fulfilled> or <not fulfilled> mark behind the sentence, indicating whether the requirement is fulfilled or not. If there is <not fulfilled> mark behind the sentence, you must modify your answer again to fulfill the requirement, by appending a new ANSWER and CHECKING part. Here is an example for this task: e.g. Question Answering Task Requirements: You need to do the Task Prompt for the following query and context. Ensure the response is written in the past tense. QUESTION: Who is Jack Chen? CONTEXTS: People saying that Jack Chen is a famous singer in China. ANSWER: Jack Chen is a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<not fulfilled> ANSWER: Jack Chen is a famous singer in China. CHECKING: Question Answering Task: You need to do the Task Prompt for the following query and context.<fulfilled>Ensure the response is written in the past tense.<fulfilled> Here is the information of your task: {Question Answering Instruction}